OpenStudy (anonymous):

Need help with rationalizing denominator (8√2)/(√20-√18)

OpenStudy (anonymous):

Multiply the numerator and denominator by the conjugate!!!! i.e \[\sqrt{20}+\sqrt{18}\] then the denominator will be of the form (a+b) * (a-b)=\[a^{2}+b ^{2}\] so u can rationalize it